OpenStudy (anonymous):

Simplify 15^-4 ------ 15^5

OpenStudy (anonymous):

its a fraction

OpenStudy (kinggeorge):

Fractions are almost the same thing as multiplying. Instead, \[\Large {a^b \over a^c}=a^{b-c}\]Using ths same principles as the last problem, and this new information, what do you think it is?

OpenStudy (anonymous):

15^-11

OpenStudy (anonymous):

15^-9 is wut i ment

OpenStudy (kinggeorge):

Precisely!\[\large {15^{-4} \over 15^5}=15^{-4-5}=15^{-9}\]

OpenStudy (kinggeorge):

Are you supposed to simplify to positive exponents?

OpenStudy (anonymous):

no it ses siplify the fraction....but 15^-9 isnt an anser

OpenStudy (kinggeorge):

Is \[1\over15^9\]an answer?

OpenStudy (anonymous):

yep..but how would it go from -9 to +9

OpenStudy (kinggeorge):

There's one more rule for exponents that's very important. \[\large a^{-b}={1 \over a^b}\]This is how you would switch between positive/negative exponents.

OpenStudy (anonymous):

ok.. so when u X u add and when u / u subtract..?

OpenStudy (kinggeorge):

Exactly.

OpenStudy (anonymous):

so wut if u had to subtract but there were 2 / ansers?

OpenStudy (kinggeorge):

What do you mean?

OpenStudy (anonymous):

In which expression should the exponents be subtracted? (6^10)^-7 2^9/5^9 (-1/2)^11 X (-1/2)^-4 3^8/3^5

OpenStudy (kinggeorge):

Only in the last one. The first option involves something we haven't gone over yet here, in the second option there are different base numbers, and in the third we have to add the numbers. In the fourth option, we have the right form, and the same base in both numbers.

OpenStudy (anonymous):

so it would be the last one?

OpenStudy (kinggeorge):

yes.

OpenStudy (anonymous):

kool